OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = (x – 3)2 – 25? Identify the x–intercepts by graphing the quadratic equation

OpenStudy (anonymous):

to find y intercept, lets x=0, to find x-intercep, lets y=0

OpenStudy (anonymous):

what would be the answer

OpenStudy (anonymous):

y int: (0-3)2- 25 = -16 xint : 0= (x-3)2 -25 25= (x-3)2 5= x-3 x= 8